The sum of two positive numbers is 50. what are the numbers if their product is to be the largest possible.




CHOICES:



25 and 25




20 and 30




24 and 26




28 and 22

1
Expert's answer
2021-11-03T17:59:12-0400

Let "x=" the first number. Then the second nmber will be "50-x."

"0<x<50"

Their product will be


"f(x)=x(50-x)=-x^2+50x"

The quadratic function "f(x)" has the absolute maximum at "x=-\dfrac{50}{2(-1)}=25."

"50-25=25."


25 and 25